OpenStudy (anonymous):

In ΔABC, which trigonometric ratio equals 32? sin B tan B tan C sin C

OpenStudy (jdoe0001):

check the sides for triangle ABC and use your SOH CAH TOA \(sin(\theta)=\cfrac{opposite}{hypotenuse} \qquad \qquad % cosine cos(\theta)=\cfrac{adjacent}{hypotenuse} \\ \quad \\ % tangent tan(\theta)=\cfrac{opposite}{adjacent}\) see which one gives you 32

OpenStudy (anonymous):

show us the triangle